# “Set It Up!”: Functional Object Arrangement with Compositional Generative Models

### v1

@@ Abstract
This paper studies the challenge of developing robots capable of understanding under-specified instructions for creating functional object arrangements, such as “set up a dining table for two”; previous arrangement approaches have focused on much more explicit instructions, such as “put object A on the table.” We introduce a framework, SetItUp, for learning to interpret under-specified instructions. SetItUp takes a small number of training examples and a human-crafted program sketch to uncover arrangement rules for specific scene types. By leveraging an intermediate graph-like representation of abstract spatial relationships among objects, SetItUp decomposes the arrangement problem into two subproblems: i) learning the arrangement patterns from limited data and ii) grounding these abstract relationships into object poses. SetItUp leverages large language models (LLMs) to propose the abstract spatial relationships among objects in novel scenes as the constraints to be satisfied; then, it composes a library of diffusion models associated with these abstract relationships to find object poses that satisfy the constraints. We validate our framework on a dataset comprising study desks, dining tables, and coffee tables, with the results showing superior performance in generating physically plausible, functional, and aesthetically pleasing object arrangements compared to existing models.
@@ I Introduction
While a large body of work has tackled the problem of generating object arrangements at a table-top or even at a house scale, most of it focused on grounding spatial relationships and making robot plans under explicit and unambiguous instructions, e.g., “put the plate next to the fork” [3, 5, 8, 19, 20, 23, 24, 34, 36]. By contrast, in this paper, we consider the task of generating object arrangements based on under-specified descriptions, such as “tidy up the study table,” “set a Chinese dinner table for two,” and “make space on a coffee table for a chess game.” Given such ambiguous instructions, our goal is to generate configurations of objects that are physically feasible, functional, aesthetic, and aligned with user preferences. We call this task orderly functional object arrangement (OFOA), and it presents three challenges.
@@
First, unlike many purely spatial relationships that have large-scale annotations [15] or can be synthetically generated according to human-written rules, global scene annotations for orderly functional object arrangements are usually scarce. This data scarcity issue becomes more severe when considering the learning of subjective preferences of a particular user. Second, not only is there very little data available, but there is also a need to make arrangement plans for a wide variety of objects, many of which might be unseen during training. The criteria for successful orderly functional object arrangement are not one-size-fits-all; they must apply to different objects and settings, leading to numerous acceptable arrangements, posing a significant challenge in generalization. Finally, pinpointing a universal measure or rules for the commonsense object arrangement is challenging due to its multifaceted nature.
@@
To tackle the under-specified and multifaceted nature of orderly functional object arrangements, our first contribution is a novel task formulation and the corresponding evaluation metrics. For a given scene type (e.g., study desks or dining tables), we formalize the task as generating the poses of objects given their category names and shapes, based on a small set of examples (we use 5 examples per scene type in this paper). These training examples include paired instructions and reference object arrangements, which can be easily collected by users. Furthermore, to enable strong and controllable generalization to unseen instructions and objects, for each scene type (e.g., dining tables), we provide a short program sketch as a hierarchical generative model specification. It essentially decomposes the arrangement task into simpler subproblems as a global “guideline” for machine learning algorithms. Finally, our benchmark also comes with a collection of rule-based metrics and human experiment rubrics for holistically evaluating different solutions in terms of their physical feasibility, functionality, and other aesthetic aspects.
@@
Our second contribution is a novel hierarchically generative approach, as illustrated in Figure 1a. Its key idea is to use a library of abstract spatial relationships, e.g., left-of, horizontally-aligned, evenly-spaced, as its intermediate prediction task. This breaks down the task of predicting object poses into two steps: generating a set of object relationships that should be satisfied in the final arrangement, and generating concrete object poses that comply with these relationships. In our system SetItUp, we leverage large language models (LLMs) for the first prediction task and use a library of compositional generative models to ground the relationships into poses.
@@ II Problem Formulation
We frame orderly functional object arrangement (OFOA), using a novel few-shot learning paradigm. Specifically, each object arrangement task family is a tuple $\langle{\mathcal{D}},\textit{desc},\mathcal{O},\textit{sketch}\rangle$, where ${\mathcal{D}}$ represents a small dataset of examples corresponding to a specific scene type and user preferences. Each training example $d\in{\mathcal{D}}$ is a tuple $\langle\textit{desc}_{d},\mathcal{O}_{d},{\mathcal{P}}_{d}\rangle$. Here, $\textit{desc}_{d}$ denotes a natural language instruction for a specific task in that family (e.g., “set up a dining table for two”) that implies a preferred object arrangement. $\mathcal{O}_{d}$ and ${\mathcal{P}}_{d}$ encode a desired scene configuration that fulfills the instruction $\textit{desc}_{d}$. Specifically, $\mathcal{O}_{d}=\{o_{0},o_{1},\ldots,o_{N-1}\}$ denotes a collection of input objects categorized by two static properties: types ($n_{i}$) and shapes ($g_{i}$, represented as 2D bounding boxes in this paper). The desired “output” arrangement ${\mathcal{P}}_{d}=\{p_{0},p_{1},\ldots,p_{N-1}\}$ is represented by a planar pose for each object, expressed isn coordinates and orientation $(x,y,\theta)$ [^1] This basic problem formulation can be extended to 3D shapes and poses.. Since we only focus on tabletop arrangement tasks in this paper, all poses are represented in a canonical table frame. We need very few (on the order of five) examples, making it easy for them to be collected by users.
@@
To enable strong and controllable generalization to an unseen set of objects and instructions, we introduce a program sketch for each task family. This sketch defines a hierarchical generative model for generating object arrangements for a particular family of tasks. Essentially, it decomposes the problem of generating the full arrangement into a hierarchy of smaller sub-problems. Each sub-problem has a simpler input-output specification, such as categorizing object types and generating the arrangement of a subset of the object. In this paper, we focus on leveraging human-defined sketch descriptions for three domains: study desks, coffee tables, and dining tables. They were written in a Python-like syntax and could conceivably be generated from free-form language instructions from humans as in Zelikman et al. [35].
@@ III Set It Up
To address the challenges of data-efficient learning and robust generalization in orderly functional object arrangement (OFOA) tasks, we introduce a novel hierarchical generative framework, namely SetItUp. The key idea of SetItUp is to use a library of abstract spatial relationships as an intermediate representation for solving the arrangement task. This library is composed of basic spatial relationships such as left-of and horizontally-aligned, which can be unambiguously defined by considering the geometric features of objects. Based on this library, we can decompose the problem of learning to generate object poses given their categories, shapes, and the task instructions into two subproblems: learning to generate the set of spatial relationships that encodes the arrangement plan, and learning to generate object poses based on a set of specified spatial relationships.
Figure 2 illustrates the overall framework. It combines large language models (LLMs) as a repository of commonsense knowledge and as a strong few-shot learner for generating abstract descriptions of the arrangement, and compositional diffusion models as powerful generative models to ground these spatial relationships into object poses. Given the task input tuple $\langle{\mathcal{D}},\textit{desc},{\mathcal{O}},\textit{sketch}\rangle$, we first convert all training examples in ${\mathcal{D}}$ into a language description based on the abstract spatial relationship library. Next, we prompt an LLM, based on the task inputs, to generate a set of abstract spatial relationships among test objects. This gives us a factor graph-like representation of the desired arrangement. Finally, we leverage a composition of the diffusion-based generative models associated with each abstract relationship to predict the output object poses associated with each object, which comply with the factor graph description generated by the LLM.
In the following sections, we will first present the library of abstract spatial relationships (Section III-A), where each relationship is associated with a simple geometric-rule-based classifier and a neural network-based generative model. Following that, we will introduce a compositional diffusion model inference algorithm capable of generating object poses based on the factor graph specification of object relationships. Finally, in Section III-C, we specify our strategy for prompting an LLM based on the training examples, test inputs, and scene-type sketches to propose the abstract spatial relationships.
@@
III-A The Abstract Spatial Relationship Library
@@
Ground graph representation of object arrangements. Given the library of abstract spatial relationships ${\mathcal{R}}$, we can encode a desired spatial arrangement of an object set ${\mathcal{O}}$ as a graph of ground spatial relations, ${\mathcal{G}}=\{r_{i}(o^{i}_{1},\ldots,o^{i}_{k_{i}})\}_{i}$ where each $r_{i}\in\mathcal{R}$ is a relationship (such as horizontally-aligned), $k_{i}$ is the arity of $r_{i}$, and $o^{i}_{1},\ldots,o^{i}_{k_{i}}$ are elements of $\mathcal{O}$. The objective is to produce a set of poses ${\mathcal{P}}=\{p_{i}\}$, such that, for all elements $(R,(o_{1},\ldots,o_{k_{R}}))\in\mathcal{G}$, $h_{R}(g_{1},\ldots g_{k_{R}},p_{1},\ldots,p_{k_{R}})$ is true, where $g_{i}$ is the corresponding shape of $o_{i}$.
@@
Classifiers for abstract spatial relationships. Since all relationships used in our examples are unambiguously defined based on simple geometric transformations (e.g., by comparing the 2D coordinate of objects in a canonical table frame), we use a small set of rules to construct the classifier function $h_{R}$. We include the details of the rules in the appendix A-C. They could instead be learned in conjunction with the generative models for relationships that are defined through examples only.
@@
Generative models for abstract spatial relationships. One straightforward approach to the overall problem would be to hand-specify $f_{R}$ for each relationship type and use standard non-linear optimization methods to search for ${\mathcal{P}}$. There are two substantial difficulties with this approach. First, we may want to extend to relationship types for which we do not know an analytical form for $f$, and so would want to acquire it via learning. Second, the optimization problem for sampling assignments to $P$ given a ground graph representation $({\mathcal{P}},{\mathcal{G}})$ is highly non-convex and hence very difficult, and the standard method would typically require a great deal of tuning (e.g., of the steepness of the objective near the constraint boundary).
@@ III-B Pose Generation via Compositional Diffusion Models
III-C Abstract Relationship Generation via Rule Induction
@@
Figure 4: Abstract relationship generation through rule induction involves two phases. Initially, in the schema induction phase, we employ an LLM to create a “setup” rule schema from a few training examples and a high-level task-family sketch defined by humans. Then, with a new instruction and the list of test objects, the LLM applies this schema to propose the final set of abstract spatial relationships as a ground graph.
Now that we have a method for going from a ground graph description into object poses, we develop a method for generating graph descriptions ${\mathcal{G}}$ using a large language model. As illustrated in Figure 4, we use the LLM in two ways. First, at training time, we learn about the task family by taking a small set of training examples, ${\mathcal{D}}$, and an abstract sketch of desired arrangements as input and producing a set of task family rule schemas. Then, at inference time, we take a new set of objects ${\mathcal{O}}$ and the task instruction desc, together with the rule schemas for this task family, to generate the ground relationship graph.
@@
Figure 5: Example process of using LLM to instantiate a program sketch. Sub-figure (a) presents an example of the initial program sketch. We provide this program sketch, along with five training instances, to the LLM. The LLM then instantiates it by adding documentation and summarization of the common patterns, as illustrated in (b). With this partially instantiated code, the LLM is tasked to interpret and propose the object relationships to new instructions and objects. An example of an instantiated function during inference time is depicted in (c).
@@
We first induce a set of rule schema from the small training set ${\mathcal{D}}$ and the high-level task-family program sketch. To encode the pose information in training examples ${\mathcal{D}}$, for each data point in the training set $(\textit{desc}_{d},{\mathcal{O}}_{d},{\mathcal{P}}_{d})\in{\mathcal{D}}$, we compute the set of primitive relationships that hold by applying the classifier $h_{R}$ for each abstract relationship type $R\in{\mathcal{R}}$, to each subset of objects in ${\mathcal{O}}_{d}$. We use a string encoding of all active relationships to describe each example scene. Since the sketch guides the LLM to decompose the relationship prediction task into sub-problems, at this stage we only generate rule schemas for individual subroutines in the sketch.
@@
Sketch. Our relationship generation procedure adopts a hierarchical program synthesis framework, similar to the approach described in Zelikman et al. [35]. It is based on a human-defined, task-family-specific sketch, with an example shown in Figure 5. Such high-level sketches decompose the generation task into several stages and four types of subroutines: 1) extracting information from the instruction (e.g., number of diners), 2) categorizing objects into groups (e.g., finding all input devices to a computer), 3) generating arrangements for a given set of objects (e.g., all input devices), and 4) generating arrangements for an object set based on the generated placements of another object set. These subroutines usually require only a single “step” of reasoning, such as directly extracting a number from the instruction or generating object arrangements for a smaller subset of the objects. In this work, we represent the sketch in Python-like syntax, but in future work, it would be important, like in Zelikman et al. [35], to accept sketches in natural language and convert them into Python or a simpler language such as their Parsel. In practice, we manually decompose the prediction task into 7-10 smaller sub-problems; we provide the sketches for the three task families we study in this work in the appendix A-B. To add a new task family, such as setting up for a Scrabble game, it is only necessary to provide a sketch and a few example set-ups. Furthermore, it is crucial to note that this high-level sketch does not contain concrete implementations for any subroutines; we will prompt the LLM, given the training examples, to solve each subproblem.
@@
Rule schema induction. Given the sketch and training examples, we ask the LLM to write documentation for each subroutine that summarises the rules associated with it, such as instructions for identifying the key object in a study desk setup task. Note that the LLM may opt to write either documentation or a partially instantiated function. The textual version of the training examples, described in terms of their spatial relations, is input to the LLM as well, and the prompt queries the LLM to summarize the patterns in the training examples into the documentation. Examples of a sketch and a resulting LLM-generated rule schema are illustrated in Figure 5a and b.
@@
Schema instantiation and execution. At performance time, the system is given a new set of objects ${\mathcal{O}}_{d}$ and task description descr. We provide this information, together with the task-family rule schemas, to the LLM and ask it to generate the final set of abstract relationships. The LLM is prompted to read the test instructions desc and the object names in ${\mathcal{O}}$, and fill in scene-specific details for each rule schema. A fully instantiated version of the schema extract_main_devices is shown in Figure 5c: the LLM generates a list of key objects for the new scene and incorporates it into a return statement. To generate the final set of relationships among objects, we simply provide the fully instantiated setup rules, which now include the scene-specific documentation and implementation, as well as the task description to the LLM and prompt it to generate the list of relationships.

### v3

@@ Abstract
This paper studies the challenge of developing robots capable of understanding under-specified instructions for creating functional object arrangements, such as “set up a dining table for two”; previous arrangement approaches have focused on much more explicit instructions, such as “put object A on the table.” We introduce a framework, SetItUp, for learning to interpret under-specified instructions. SetItUp takes a small number of training examples and a human-crafted program sketch to uncover arrangement rules for specific scene types. By leveraging an intermediate graph-like representation of abstract spatial relationships among objects, SetItUp decomposes the arrangement problem into two subproblems: i) learning the arrangement patterns from limited data and ii) grounding these abstract relationships into object poses. SetItUp leverages large language models (LLMs) to propose the abstract spatial relationships among objects in novel scenes as the constraints to be satisfied; then, it composes a library of diffusion models associated with these abstract relationships to find object poses that satisfy the constraints. We validate our framework on a dataset comprising study desks, dining tables, and coffee tables, with the results showing superior performance in generating physically plausible, functional, and aesthetically pleasing object arrangements compared to existing models.  [^1] Project page: https://setitup-rss.github.io/
@@ I Introduction
While a large body of work has tackled the problem of generating object arrangements at a table-top or even at a house scale, most of it focused on grounding spatial relationships and making robot plans under explicit and unambiguous instructions, e.g., “put the plate next to the fork” [3, 5, 8, 19, 20, 23, 24, 34, 36]. By contrast, in this paper, we consider the task of generating object arrangements based on under-specified descriptions, such as “tidy up the study table,” “set a Chinese dinner table for two,” and “make space on a coffee table for a chess game.” Given such ambiguous instructions, our goal is to generate configurations of objects that are physically feasible, functional, aesthetic, and aligned with user preferences. We call this task functional object arrangement (FORM), and it presents three challenges.
@@
First, unlike many purely spatial relationships that have large-scale annotations [15] or can be synthetically generated according to human-written rules, global scene annotations for functional object arrangements are usually scarce. This data scarcity issue becomes more severe when considering the learning of subjective preferences of a particular user. Second, not only is there very little data available, but there is also a need to make arrangement plans for a wide variety of objects, many of which might be unseen during training. The criteria for successful functional object arrangement are not one-size-fits-all; they must apply to different objects and settings, leading to numerous acceptable arrangements, posing a significant challenge in generalization. Finally, pinpointing a universal measure or rules for the commonsense object arrangement is challenging due to its multifaceted nature.
@@
To tackle the under-specified and multifaceted nature of functional object arrangements, our first contribution is a novel task formulation and the corresponding evaluation metrics. For a given scene type, e.g., study desks or dining tables, we formalize the task as generating the poses of objects given their category names and shapes, based on a small set of examples. We used 5 examples per scene type in our experiments. These training examples include paired instructions and reference object arrangements, which can be easily collected by users. Furthermore, to enable strong and controllable generalization to unseen instructions and objects, for each scene type, we provide a short program sketch written in Python-like domain-specific language as a hierarchical generative model specification. The sketch provides a list of function names, signatures, and descriptions, but does not include any implementation. Essentially, it decomposes the arrangement task into simpler subproblems as a global “guideline” for machine learning algorithms. Finally, our benchmark also comes with a collection of rule-based metrics and human experiment rubrics for holistically evaluating different solutions in terms of their physical feasibility, functionality, and other aesthetic aspects.
@@
Our second contribution is a novel hierarchically generative approach, as illustrated in Figure 1a. Its key idea is to use a library of abstract spatial relationships, e.g., left-of, horizontally-aligned, as its intermediate prediction task. This breaks down the task of predicting object poses into two steps: generating a set of object relationships that should be satisfied in the final arrangement, and generating concrete object poses that comply with these relationships. In our system SetItUp, we leverage large language models (LLMs) for the first prediction task and use a library of compositional generative models to ground the relationships into poses.
@@ II Problem Formulation
We frame functional object arrangement (FORM), using a novel few-shot learning paradigm. Specifically, each object arrangement task family is a tuple $\langle{\mathcal{D}},\textit{desc},\mathcal{O},\textit{sketch}\rangle$, where ${\mathcal{D}}$ represents a small dataset of examples corresponding to a specific scene type and user preferences. Each training example $d\in{\mathcal{D}}$ is a tuple $\langle\textit{desc}_{d},\mathcal{O}_{d},{\mathcal{P}}_{d}\rangle$. Here, $\textit{desc}_{d}$ denotes a natural language instruction for a specific task in that family (e.g., “set up a dining table for two”) that implies a preferred object arrangement. $\mathcal{O}_{d}$ and ${\mathcal{P}}_{d}$ encode a desired scene configuration that fulfills the instruction $\textit{desc}_{d}$. Specifically, $\mathcal{O}_{d}=\{o_{0},o_{1},\ldots,o_{N-1}\}$ denotes a collection of input objects categorized by two static properties: types ($n_{i}$) and shapes ($g_{i}$, represented as 2D bounding boxes in this paper). The desired “output” arrangement ${\mathcal{P}}_{d}=\{p_{0},p_{1},\ldots,p_{N-1}\}$ is represented by a planar pose for each object, expressed isn coordinates and orientation $(x,y,\theta)$ [^2] This basic problem formulation can be extended to 3D shapes and poses.. Since we only focus on tabletop arrangement tasks in this paper, all poses are represented in a canonical table frame. We need very few (on the order of five) examples, making it easy for them to be collected by users.
@@
To enable strong and controllable generalization to an unseen set of objects and instructions, we introduce a program sketch for each task family. This sketch defines a hierarchical generative model for generating object arrangements for a particular family of tasks. It uses a Python-like syntax to outline the task-solving procedures with only function names, signatures, and descriptions, but no implementations. Essentially, it decomposes the problem of generating the full arrangement into a hierarchy of smaller sub-problems. Each sub-problem has a simpler input-output specification, such as categorizing object types and generating the arrangement of a subset of the object. In this paper, we focus on leveraging human-defined sketch descriptions for three domains: study desks, coffee tables, and dining tables. Moving forward, the sketches could conceivably be generated from free-form language instructions from humans as in Zelikman et al. [35].
@@ III Set It Up
To address the challenges of data-efficient learning and robust generalization in functional object arrangement (FORM) tasks, we introduce a novel hierarchical generative framework, namely SetItUp. The key idea of SetItUp is to use a library of abstract spatial relationships to construct a grounding graph, which serves as an intermediate representation for solving the arrangement task. This library is composed of basic spatial relationships such as left-of and horizontally-aligned, which can be unambiguously defined by considering the geometric features of objects. Based on this library, we can decompose the problem of learning to generate object poses given their categories, shapes, and the task instructions into two subproblems: learning to generate the set of spatial relationships that encodes the arrangement plan, and learning to generate object poses based on a set of specified spatial relationships.
Figure 2 illustrates the overall framework. Central to our framework is the use of abstract spatial relationships as the intermediate representation. This high-level abstraction over object poses naturally breaks down FORM tasks into two generative sub-tasks. The first sub-task involves a LLM-based symbolic generative model that generates the functional abstract relationships for the test objects. The second sub-task involves a diffusion-based generative model that takes these abstract relationships as input and propose the corresponding object poses. Essentially, SetItUp combines large language models (LLMs) as a repository of commonsense knowledge and as a strong few-shot learner for generating abstract descriptions of the arrangement, and compositional diffusion models as powerful generative models to ground these spatial relationships into object poses. Given the task input tuple $\langle{\mathcal{D}},\textit{desc},{\mathcal{O}},\textit{sketch}\rangle$, we first convert all training examples in ${\mathcal{D}}$ into a language description based on the abstract spatial relationship library. Next, we prompt an LLM, based on the task inputs, to generate a set of abstract spatial relationships among test objects. This gives us a factor graph-like representation of the desired arrangement. Finally, we leverage a composition of the diffusion-based generative models associated with each abstract relationship to predict the output object poses associated with each object, which comply with the factor graph description generated by the LLM.
In the following sections, we will first present the library of abstract spatial relationships (Section III-A), where each relationship is associated with a simple geometric-rule-based classifier and a neural network-based generative model. Following that, we will introduce a compositional diffusion model inference algorithm capable of generating object poses based on the factor graph specification of object relationships. Finally, in Section III-C, we specify our strategy for prompting an LLM based on the training examples, test inputs, and task-family sketches to propose the abstract spatial relationships.
@@
III-A The Spatial Relationship Library
@@
Grounding graph. Given the library of abstract spatial relationships ${\mathcal{R}}$, we can encode a desired spatial arrangement of an object set ${\mathcal{O}}$ as a graph of ground spatial relations, ${\mathcal{G}}=\{r_{i}(o^{i}_{1},\ldots,o^{i}_{k_{i}})\}_{i}$ where each $r_{i}\in\mathcal{R}$ is a relationship (such as horizontally-aligned), $k_{i}$ is the arity of $r_{i}$, and $o^{i}_{1},\ldots,o^{i}_{k_{i}}$ are elements of $\mathcal{O}$. The objective is to produce a set of poses ${\mathcal{P}}=\{p_{i}\}$, such that, for all elements $(R,(o_{1},\ldots,o_{k_{R}}))\in\mathcal{G}$, $h_{R}(g_{1},\ldots g_{k_{R}},p_{1},\ldots,p_{k_{R}})$ is true, where $g_{i}$ is the corresponding shape of $o_{i}$.
@@
Classifying spatial relationships Since all relationships used in our examples are unambiguously defined based on simple geometric transformations (e.g., by comparing the 2D coordinate of objects in a canonical table frame), we use a small set of rules to construct the classifier function $h_{R}$. We include the details of the rules in the appendix A-C. They could instead be learned in conjunction with the generative models for relationships that are defined through examples only.
@@
Generating spatial relationships. One straightforward approach to the overall problem would be to hand-specify $f_{R}$ for each relationship type and use standard non-linear optimization methods to search for ${\mathcal{P}}$. There are two substantial difficulties with this approach. First, we may want to extend to relationship types for which we do not know an analytical form for $f$, and so would want to acquire it via learning. Second, the optimization problem for sampling assignments to $P$ given a ground graph representation $({\mathcal{P}},{\mathcal{G}})$ is highly non-convex and hence very difficult, and the standard method would typically require a great deal of tuning (e.g., of the steepness of the objective near the constraint boundary).
@@ III-B Pose Generation via Compositional Diffusion Models
III-C Abstract Relationship Generation via Program Induction
@@
Figure 4: Abstract relationship generation through rule induction involves two phases. Initially, in the program induction phase, we employ an LLM to create a “setup” rule-based program from a few training examples and a high-level task-family sketch defined by humans. This program contains rules and patterns for various subproblems, but it has unbound variables (i.e., the actual objects and instructions are not specified yet). In the second phase, with a new instruction and a list of test objects, the LLM binds these variables to the induced program to create an executable Python program. This executable Python program is then used to generate the final set of abstract spatial relationships as a ground graph.
Now that we have a module for going from a ground graph description into object poses, we further generate graph descriptions ${\mathcal{G}}$ using a large language model in this section. As shown in Figure 4, our approach involves a two-stage generative process using the LLM. At the training stage, we synthesize a task family rule-based program that captures patterns for object arrangements. This program comprises detailed comments and code templates with unbound variables, as the specific objects and instructions are not yet defined. The synthesis process uses a small set of training examples, ${\mathcal{D}}$, to instantiate a provided sketch. During the inference stage, given a new set of objects ${\mathcal{O}}$ and a task instruction desc, we bind the actual variables to the induced program, producing an executable Python program. This script is then executed to generate the ground relationship graph.
@@
Figure 5: Example process of using an LLM to instantiate a program sketch. Sub-figure (a) presents an example of the initial program sketch. We provide this program sketch, along with five training instances, to the LLM. The LLM then creates a rule-based program, summarizing the common patterns in the form of code comments and/or templates, but with unbound variables, as illustrated in (b). Finally, given new objects and instructions, the LLM binds these variables to the induced program and generates an executable Python program. This program is then used to generate the object grounding graph. An example of an executable Python program with variable bindings during inference time is depicted in (c).
@@
Recall that the training examples only record object poses, these numerical values cannot be effectively used by the LLM to infer abstract relationship patterns. Therefore, we need to “translate” the pose information in training examples ${\mathcal{D}}$ into a set of active abstract relationships. To do this, for each data point in the training set $(\textit{desc}_{d},{\mathcal{O}}_{d},{\mathcal{P}}_{d})\in{\mathcal{D}}$, we compute the set of primitive relationships that hold by applying the classifier $h_{R}$ for each abstract relationship type $R\in{\mathcal{R}}$, to each subset of objects in ${\mathcal{O}}_{d}$. We use a string encoding of all active relationships to describe each example scene.
@@
Program Sketch. Our relationship generation procedure adopts a hierarchical program synthesis framework. It is based on a human-defined, task-family-specific sketch, with an example shown in Figure 5. Each sketch includes several functions with meaningful names, signatures, and descriptions, organized in a sequence that outlines the steps to solve the complex task. The sketch is crucial for generating abstract relationships. It acts as a guide that helps the LLM generate object relationships step-by-step. With this strong guidance, the LLM can create task-family-specific rule-based program from just five examples. Instead of directly regressing the final relationship proposals on these five training examples, we use the examples mainly to instantiate the subroutines in the sketch and to write detailed comments. Such high-level sketches decompose the generation task into four types of subroutines: 1) extracting the task relevant information from the instruction (e.g., number of diners), 2) categorizing objects into groups (e.g., finding all input devices to a computer), 3) generating arrangements for the objects within each group(e.g., all input devices), and 4) generating arrangements for objects among groups. These subroutines usually require only a single “step” of reasoning, such as directly extracting a number from the instruction or generating object arrangements for a smaller subset of the objects. In this work, we represent the sketch in Python-like syntax, but in future work, it would be important, like in Zelikman et al. [35], to accept sketches in natural language and convert them into Python or a simpler language such as their Parsel. In practice, we follow the four subroutines and manually decompose the prediction task into 7-10 functions; we provide the sketches for the three task families we study in this work in the appendix A-B. To add a new task family, such as setting up for a Scrabble game, it is only necessary to provide a sketch and a few example set-ups. Furthermore, it is crucial to note that this high-level sketch does not contain concrete implementations for any subroutines; we will prompt the LLM, given the training examples, to solve each subproblem.
@@
Program Induction. Given the sketch and training examples, we query the LLM to generate a rule-based program. This program comprising rules and patterns for each subproblem, but with unbound variables (i.e., the actual objects and instructions are not yet specified). Specifically, the program encapsulates the rules associated with each function in the form of code comments and templates. For example, it may include instructions for identifying the key object in a study desk setup task. During the rule induction stage, we input the textual descriptions of the training examples, which are detailed in terms of their spatial relations, to the LLM. We then prompt the LLM to summarize the patterns in these examples into either comments or code templates with unbound variables. After the rule induction stage, we obtain a program with comments that summarizes the patterns derived from the training examples. Examples of a sketch and a resulting LLM-generated program are illustrated in Figure 5a and b.
@@
Variable Binding and Program Execution. At performance time, given a new set of objects ${\mathcal{O}}_{d}$, task description descr, we first use the LLM to bind these variables to the induced program. This process incorporates the contextual information to generate an executable Python program. Then, we execute the program to return the set of functional abstract spatial relationships. Specifically, we prompt the LLM to read the test instructions desc and the object names in ${\mathcal{O}}$, and fill in scene-specific details for each program. An example of an executable program with variable bindings, extract_main_devices, is shown in Figure 5c: the LLM generates a list of key objects for the new scene and incorporates it into a return statement. To generate the final set of relationships among objects, we simply execute this program with variable bindings, which now include the scene-specific comments and implementation, as well as the task description to the LLM, to generate the list of relationships.